Find value of 'a' when y=(4a-6)x and y= 1/5x
Answer by It is costly(175)

find a; given;y=(4a-6)x and y=1/5x
y=4ax-6x
1/5x=4ax-6x
(5x)(1/5x)=(5x)(4ax)-(5x)(6x)
1=20ax^2-30x^2
30x^2+1=20ax^2-30x^2+30x^2
30x^2+1=20ax^2
(30x^2+1)/(20x^2)=(20ax^2)/(20x^2)
(30x^2+1)/(20x^2)=a